Sportly, Inc. completed Job No. B14 during 2019. The job cost sheet listed the following: Direct materials used $100,000 Direct labor $80,000 Manufacturing overhead applied $70,000 Units produced 5,000 units Units on hand 2,000 units How much is the cost of the finished goods sold from this job?
Answer #1

Total cost of the job = Direct Materials 100,000 + Direct Labor 80,000 + Manufacturing Overhead applied 70,000

= 250,000

Cost per unit produced = 250,000/5,000 = 50 per unit

Units sold = Units produced - Units on hand

= 5,000 - 2,000 = 3,000 units

Cost of finished goods sold = 50 * 3,000

= $150,000